OpenStudy (anonymous):

Can some one show me all the steps to evaluate the Integral of (tan^3xsec^3x) dx

OpenStudy (anonymous):

\[\int\limits(\tan^3xsex^3x)dx \] Thats the problem I have to integrate...

hartnn (hartnn):

did you try u=sec x ?

hartnn (hartnn):

\(\int \tan x\sec x\sec^2x(\sec^2x-1)dx\)

hartnn (hartnn):

then put u=sec x du ....?

OpenStudy (anonymous):

du= sec x tan x?

hartnn (hartnn):

almost! du =sec x tan x dx

OpenStudy (anonymous):

Haha lol yeah I always forget the dx :P

hartnn (hartnn):

\(\int u^2(u^2-1)du\) pretty easier to integrate this.

OpenStudy (anonymous):

Okay and you replaced tan^2 with (sec^2x-1)?

hartnn (hartnn):

yes...

hartnn (hartnn):

doubts ? any kind ?

OpenStudy (anonymous):

So its U^3/3(U^3/3-1)du?

OpenStudy (anonymous):

Then I just put back u, which is sec x

hartnn (hartnn):

what! no... :( \(\int u^4-u^2 dx\)

hartnn (hartnn):

*du

OpenStudy (anonymous):

Ohh oops XD I didnt know if you were supposed to do the algebra there or not :P

hartnn (hartnn):

now integrate

OpenStudy (anonymous):

Okay so now its u^5/5-u^3/3?

OpenStudy (anonymous):

Then replace u with sec x?

hartnn (hartnn):

now, yes!

OpenStudy (anonymous):

So sec^5/5 - Sec^3/3

hartnn (hartnn):

+c and correct.

OpenStudy (anonymous):

Thanks! :D

hartnn (hartnn):

welcome ^_^